A repaarcher measures ob satisfaction among man od, single and dnorcoc cmployces to dctcrmine whether manta status can nfluencc ob satisfaction. Based on the following description in APA format, statc the valuc fork, and n. ก็ one-way analysis ot variance showed that Job satistaction did not vary by marital status,凡3, 40)-1.61, ρ > 0.05. eBook 18 points I PrevauE row Prveer Stat3 12.E c23. My Notes Ask The folloing is an incomplete table summarizing the results of a study of the variance of life satisfaction scores among unemployed, retired, part time, and full time employees, of Between groups wilhin groups (error Total 16 35 (a) Completa the F-tabla. Rund your values far mean squares and F to two dacimal placea.) of v Batween groups within groups emor Total 36 128 (b] Compute omega-squared (. (Round your answer to two decimal places.) (c) Is the decision to retain or reject the null hypothes s7 (Assume lpha equal to0.05.] O Retain the null hypothasis. Reject the ull hypothesis You may need to use the appropriate table in Appandix C to aswer this question.
Answer #1

For the first question, we are given F(3,40)=1.61.

We know that Fsim F (k-1,N-k)

Rightarrowk-1=3

Rightarrow k=3+1=4

Also, N-k=40

Rightarrow N=40+k=40+4=44

Also, we know Vnk

Rightarrow n=N/k=44/4=11

Thus, k = 4, N = 44, n = 11.

Q2)

tha sais fashion lavel o emplayed loyed retired, fu tirna & paut tima em → 서 classes k- ups l classes Nous, tablo is obtained as follouss SourassS Betwean roups df groups MSxdf - 3 x 16 Within 8 o 80136 Total SST 12839 INAL AN SHER TABLE Sousce8 48 8 80 3C 2.2222. 128 39 7-21 b)u SST + MS uoithin 48 32.22 128 + 2.22 0:31746275530.32 Cnitical Valua- F(8,36,o 05) 2 8668